When 1000 is divided by some positive integer x, the remainder is 45. What is the sum of the largest and smallest value of x
Answer by ikleyn(52835)

.
1.  The condition is equivalent to the fact that  1000-45 = 955  is divisible by  x.


2.  955 = 5*191,  and both  5  and  191  are prime numbers.


3.  Hence,  EITHER x= 5  OR  x= 191.


4.  Thus  5 + 191 = 196  is the answer.

Solved.